# Multimodal Human-Autonomous Agents Interaction Using Pre-Trained Language and Visual Foundation Models

### v1

@@ Abstract.
In this paper, we extended the method proposed in (Linus and Elmar, 2024) to enable humans to interact naturally with autonomous agents through vocal and textual conversations. Our extended method exploits the inherent capabilities of pre-trained large language models (LLMs), multimodal visual language models (VLMs), and speech recognition (SR) models to decode the high-level natural language conversations and semantic understanding of the robot’s task environment, and abstract them to the robot’s actionable commands or queries. We performed a quantitative evaluation of our framework’s natural vocal conversation understanding with participants from different racial backgrounds and English language accents. The participants interacted with the robot using both spoken and textual instructional commands. Based on the logged interaction data, our framework achieved $87.55\%$ vocal commands decoding accuracy, $86.27\%$ commands execution success, and an average latency of $0.89$ seconds from receiving the participants’ vocal chat commands to initiating the robot’s actual physical action. The video demonstrations of this paper can be found at https://linusnep.github.io/MTCC-IRoNL/.
@@ 2. Related Work
Further, L. Nwankwo et al. (Linus and Elmar, 2024) incorporated text-based interaction techniques into autonomous robots leveraging LLMs and VLMs. Nonetheless, the framework lacks complete naturalness due to the absence of a mechanism to understand vocal instructions. In this work, we build upon the foundation provided in (Linus and Elmar, 2024). Instead of relying on the accuracy of the SR model to plan the robot’s actions or depend on the text-based approach as a standalone, we propose a dual-modality approach that synergizes both the textual and vocal modalities. We leveraged the LLMs and the SR models’ abilities to maintain robustness in diverse environments. In environments where ambient noise levels may compromise the accuracy of the SR model’s vocal instructions decoding, our framework provides the flexibility to revert to the text-based interaction method. Conversely, in a quieter environment, the user can leverage the vocal modality pipeline for more natural and seamless interaction.
@@ 3. Method
In this section, we provide details on the incorporation of the vocal conversation understanding pipeline. For details about the implementation of the first four sections (the area enclosed in red dotted line in Figure 1) and how the pre-trained LLMs and VLMs are prompted to generate the actions used by the REM node, we refer the reader to (Linus and Elmar, 2024).
@@ 4.1. Preliminary Results
We invited $5$ participants (average age of $27\ (\pm 3)$ and gender distribution, $80\%$ male and $20\%$ female) with different English language accents to interact with the robot via natural vocal conversation. We logged the interaction data i.e., the SR models’ transcription of the participants’ spoken words, the LLMNode predicted labels, the true action labels, etc. We used the logged interaction data to quantitatively evaluate the performance of our framework. We defined vocal commands understanding accuracy (VCUA) metric in addition to the navigation success rate (NSR), object identification accuracy (OIA), and average response time (ART) metrics utilised in (Linus and Elmar, 2024) to assess our framework’s performance. With the VCUA, we assess how accurately the LLMNode predicts the commands based on the transcribed vocal instructions from the SRNode. We computed the accuracy as the percentage proportion of the correctly transcribed instruction to the generated instructions (from LLMNode) fed to the REM node for the actual robot’s execution.
@@
xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx
The top-left figure shows the VCUA, NSR, and OIA metrics for selected labels. We achieved $87.55\%$ VCUA and $86.27\%$ NSR, which indicates a good level of accuracy in the vocal commands decoding. In comparison to the results obtained in the textual-based method (Linus and Elmar, 2024), we observed a slight difference in the command recognition accuracy, with VCUA achieving about $12\%$ less than the textual-based CRA ($99.13\%$) as well as a $11.69\%$ reduction in the NSR. This is expected because of the ambient environmental noise and variation in the participant’s accents, which affect the vocal transcription from the SRNode, as could be seen in the confusion matrix (bottom-left of Figure 2).

### v2

@@ Abstract.
In this paper, we extended the method proposed in (Nwankwo and Rueckert, 2024) to enable humans to interact naturally with autonomous agents through vocal and textual conversations. Our extended method exploits the inherent capabilities of pre-trained large language models (LLMs), multimodal visual language models (VLMs), and speech recognition (SR) models to decode the high-level natural language conversations and semantic understanding of the robot’s task environment, and abstract them to the robot’s actionable commands or queries. We performed a quantitative evaluation of our framework’s natural vocal conversation understanding with participants from different racial backgrounds and English language accents. The participants interacted with the robot using both spoken and textual instructional commands. Based on the logged interaction data, our framework achieved $87.55\%$ vocal commands decoding accuracy, $86.27\%$ commands execution success, and an average latency of $0.89$ seconds from receiving the participants’ vocal chat commands to initiating the robot’s actual physical action. The video demonstrations of this paper can be found at https://linusnep.github.io/MTCC-IRoNL/.
@@ 2. Related Work
Further, L. Nwankwo et al. (Nwankwo and Rueckert, 2024) incorporated text-based interaction techniques into autonomous robots leveraging LLMs and VLMs. Nonetheless, the framework lacks complete naturalness due to the absence of a mechanism to understand vocal instructions. In this work, we build upon the foundation provided in (Nwankwo and Rueckert, 2024). Instead of relying on the accuracy of the SR model to plan the robot’s actions or depend on the text-based approach as a standalone, we propose a dual-modality approach that synergizes both the textual and vocal modalities. We leveraged the LLMs and the SR models’ abilities to maintain robustness in diverse environments. In environments where ambient noise levels may compromise the accuracy of the SR model’s vocal instructions decoding, our framework provides the flexibility to revert to the text-based interaction method. Conversely, in a quieter environment, the user can leverage the vocal modality pipeline for more natural and seamless interaction.
@@ 3. Method
In this section, we provide details on the incorporation of the vocal conversation understanding pipeline. For details about the implementation of the first four sections (the area enclosed in red dotted line in Figure 1) and how the pre-trained LLMs and VLMs are prompted to generate the actions used by the REM node, we refer the reader to (Nwankwo and Rueckert, 2024).
@@ 4.1. Preliminary Results
We invited $5$ participants (average age of $27\ (\pm 3)$ and gender distribution, $80\%$ male and $20\%$ female) with different English language accents to interact with the robot via natural vocal conversation. We logged the interaction data i.e., the SR models’ transcription of the participants’ spoken words, the LLMNode predicted labels, the true action labels, etc. We used the logged interaction data to quantitatively evaluate the performance of our framework. We defined vocal commands understanding accuracy (VCUA) metric in addition to the navigation success rate (NSR), object identification accuracy (OIA), and average response time (ART) metrics utilised in (Nwankwo and Rueckert, 2024) to assess our framework’s performance. With the VCUA, we assess how accurately the LLMNode predicts the commands based on the transcribed vocal instructions from the SRNode. We computed the accuracy as the percentage proportion of the correctly transcribed instruction to the generated instructions (from LLMNode) fed to the REM node for the actual robot’s execution.
@@
The top-left figure shows the VCUA, NSR, and OIA metrics for selected labels. We achieved $87.55\%$ VCUA and $86.27\%$ NSR, which indicates a good level of accuracy in the vocal commands decoding. In comparison to the results obtained in the textual-based method (Nwankwo and Rueckert, 2024), we observed a slight difference in the command recognition accuracy, with VCUA achieving about $12\%$ less than the textual-based CRA ($99.13\%$) as well as a $11.69\%$ reduction in the NSR. This is expected because of the ambient environmental noise and variation in the participant’s accents, which affect the vocal transcription from the SRNode, as could be seen in the confusion matrix (bottom-left of Figure 2).